OpenStudy (anonymous):

Brenda drives an 18 wheeler. she plans to buy her own truck. Her research indicates that the expected running costs, C, in dollars, per 100 km, are given by C(v)=0.9 +0.0016v square, where v is the speed, in kilometers per hour. Brenda's first trip will be 1500 km, round trip. she plans to pay herself $30/h. determine the speed that will minimize Brend's cost for the trip

OpenStudy (anonymous):

x(0.0016x^2+0.9)/100=30 x=121.790 Total cost of trip (1500/x)*(x(0.0016x^2+0.9/100)=15(0.0016x^2+0.9) x is the speed in km/hr, so that cost is 30 $/hr. Any speed less than that would result in lower $/hr as well as lower cost of the trip.

OpenStudy (anonymous):

The answer in the textbook says 97.9 km/hr

OpenStudy (anonymous):

The answer in the textbook says 97.9 km/hr